OpenStudy (anonymous):

Dolls cost $140 per carton and trucks cost $430 per carton. If an order comes in for a total of 100 cartons for $28,500, what was the number of cartons of dolls? (Hint, Let T = Cartons of Trucks

OpenStudy (campbell_st):

set up the 2 equations D + T = 100 (1) total number of cartons 140D + 430T = 28500 (2) total cost rewrite equation 1 T = 100 - D and substitute into equation 2 140D + 430(100 - D) = 28500 expand 140D + 43000 - 430D = 28500 collect like terms - 290D = -14500 \[D = \frac{-14500}{- 290}\] then there were 50 cartons of dolls and 50 cartons of trucks hope this helps